OpenStudy (jcmbourque19):

Quick math help please!!! What is log2 (3) + log2 (x)=3? Also how can the properties of exponents help solve logarithmic equations? I'm stumped on these two and I've been trying to figure them out for so long.

OpenStudy (raden):

use the property : loga (m) + loga (n) = loga (mn)

OpenStudy (jcmbourque19):

so that would be log2 (3x). What do I do next?

OpenStudy (raden):

yeah, then the equation becomes log2 (3x) = 3, right ?

OpenStudy (jcmbourque19):

I still don't get how to solve... really don't understand this

OpenStudy (raden):

log2 (3x) = 3 we can convert it into exponent form : 2^3 = 3x then solve for x

OpenStudy (raden):

got it ?

OpenStudy (jcmbourque19):

8=3x so x=8/3 Thank you so much!!!

OpenStudy (raden):

yes, that's right. welcome